. What will be the unit digit of the squares of the following numbers? i. 81 ii. 272 iii. 799 iv. 3853  
gyklutld7tdltld

Grade:12th pass

1 Answers

Chetan Jaiswal
2 years ago
Hello!
 
  1. 81:- (1*1 = 1), so the units digit of the square of 81 is ‘1’
  2. 272 :– (2*2 = 4), so the units digit of the square of 272 is ‘4’
  3. 799 :– (9*9 = 8), so the units digit of the square of 799 will be ‘1’
  4. 3853 :- (3*3 = 9), so the units digit of the square of 3853 will be ‘9’
Thanks!